# UniEgoMotion: A Unified Model for Egocentric Motion Reconstruction, Forecasting, and Generation

### v1

@@ 3.5 Motion Representation
To address these issues, we adopt a head-centric representation. We transform into using forward kinematics where is the global SE(3) transform of the head joint, and are those of other joints. This removes joint dependencies in the kinematic chain. Next, we derive a canonical reference frame per frame by eliminating pitch, roll, and height relative to the floor, ensuring that captures the head’s global trajectory projected onto the floor. Motion is then expressed as where the latter terms encode local canonicalized pose information. For trajectory invariance, we represent as its residual relative to the previous frame. While [92] adopts a similar canonicalization scheme, it preserves the kinematic chain, resulting in severe foot-floor penetration and floating artifacts. Our experiments validate the effectiveness of our motion representation against [92].
@@ 4.3 Results & Discussion
Egocentric Motion Reconstruction: Tab. 1 shows the results and ablation study for the egocentric reconstruction task, and Fig. 3 shows the qualitative comparison of the reconstruction baselines. UniEgoMotion’s egocentric reconstruction capabilities outperform specialized baselines – AvatarPoser[42], EgoEgo [48], and EgoAllo [92], in both reconstruction and semantic metrics. EgoAllo’s pelvis-centric motion representation struggles to accurately follow egocentric inputs, resulting in high head rotation and translation errors. As seen in Fig. 3, it shows frequent floating motion which explains its high Foot Contact error in Tab. 1. We further verify the benefits of egocentric motion representation by training UniEgoMotion with a pelvis-centric representation. We also show that transformer cross-attention is better suited for our flexible conditioning setting [72] compared to encoder-based architectures [48, 30] and specialized 1D-Unet [44]. Thanks to the explicit use of fine-grained image features, UniEgoMotion captures visible semantic cues more effectively and achieves the lowest MPJPE-* errors along with the highest motion quality and semantic similarity to the ground-truth motion.
@@ Appendix B Baselines
We compare the egocentric motion reconstruction capabilities of UniEgoMotion with task-specific prior works: EgoEgo [48], EgoAllo [92], and AvatarPoser [42]. To ensure a fair evaluation, we retrain each method on the EE4D-Motion dataset using their publicly available code. Following EgoEgo’s experimental setup, we exclude hand tracking from AvatarPoser and instead provide a constant input for hand trajectories. Both EgoEgo and AvatarPoser use head trajectories derived from motion annotations rather than from the Aria device’s SLAM system, resulting in perfect head tracking by design. Therefore, we omit their head tracking metrics from the evaluation. For EgoAllo, we evaluate the output of the motion diffusion model directly, without applying the post-processing optimization step.
@@ Appendix C Ablation on Conditioning Inputs
Appendix D Why Not Text Conditioning
Appendix E Training Details
Appendix F Motion Representation
@@ Appendix G EE4D-Motion Dataset
To bridge this gap, we process the large-scale EgoExo4D dataset [23] to generate pseudo-ground-truth 3D motion data. We refer to this processed dataset as EE4D-Motion, which consists of 110+ hours of time-synchronized 3D motion data and egocentric videos, alongside other EgoExo4D annotations. This dataset serves as an extensive benchmark for multimodal motion research.
@@
Filtering & Quality Control: We filter out segments with excessive jitter caused by erroneous device trajectories, suboptimal off-the-shelf model predictions, or severe occlusions across all exo views.

### v2

@@ 3.5 Motion Representation
To address these issues, we adopt a head-centric representation. We transform into using forward kinematics where is the global SE(3) transform of the head joint, and are those of other joints. This removes joint dependencies in the kinematic chain. Next, we derive a canonical reference frame per frame by eliminating pitch, roll, and height relative to the floor, ensuring that captures the head’s global trajectory projected onto the floor. Motion is then expressed as where the latter terms encode local canonicalized pose information. For trajectory invariance, we represent as its residual relative to the previous frame. For more details, please refer to the supplementary material. While [92] adopts a similar canonicalization scheme, it preserves the kinematic chain, resulting in severe foot-floor penetration and floating artifacts. Our experiments validate the effectiveness of our motion representation against [92].
@@ 4.3 Results & Discussion
Egocentric Motion Reconstruction: Tab. 1 shows the results and ablation study for the egocentric reconstruction task, and Fig. 3 shows the qualitative comparison of the reconstruction baselines. UniEgoMotion’s egocentric reconstruction capabilities outperform specialized baselines – AvatarPoser[42], EgoEgo [48], and EgoAllo [92], in both reconstruction and semantic metrics. EgoAllo’s motion representation with a kinematic chain and local joint angles, often struggles to reconstruct motion accurately grounded on the floor. As a result, it shows frequent foot sliding and floating motion, leading to its high Foot Contact error in Tab. 1. We further verify the benefits of egocentric motion representation by training UniEgoMotion with a pelvis-centric representation. We also show that transformer cross-attention is better suited for our flexible conditioning setting [72] compared to encoder-based architectures [48, 30] and specialized 1D-Unet [44]. Thanks to the explicit use of fine-grained image features, UniEgoMotion captures visible semantic cues more effectively and achieves the lowest MPJPE-* errors along with the highest motion quality and semantic similarity to the ground-truth motion.
@@ Appendix B Baselines
We compare the egocentric motion reconstruction capabilities of UniEgoMotion with task-specific prior works: EgoEgo [48], EgoAllo [92], and AvatarPoser [42]. To ensure a fair evaluation, we retrain each method on the EE4D-Motion dataset using their publicly available code. Following EgoEgo’s experimental setup, we exclude hand tracking from AvatarPoser and instead provide a constant input for hand trajectories. Both EgoEgo and AvatarPoser use as inputs the head trajectories derived from motion annotations rather than from the Aria device’s SLAM system, resulting in perfect head tracking by design. Similarly, EgoAllo also uses a fixed transformation between head and device trajectory. Therefore, we omit their head tracking metrics from the evaluation. For EgoAllo, we evaluate the output of the motion diffusion model directly, without applying the post-processing optimization step.
@@ Appendix C Ablation on Conditioning Inputs
Appendix D Motion Representation
@@
Appendix E Why Not Text Conditioning
Appendix F Training Details
@@ Appendix G EE4D-Motion Dataset
To bridge this gap, we process the large-scale EgoExo4D dataset [23] to generate pseudo-ground-truth 3D motion data. We refer to this processed dataset as EE4D-Motion, which consists of 208 hours of time-synchronized 3D motion data and egocentric videos, alongside other EgoExo4D annotations. This dataset serves as an extensive benchmark for multimodal motion research.
@@
Filtering & Quality Control: We filter out segments with excessive jitter caused by erroneous device trajectories, suboptimal off-the-shelf model predictions, or severe occlusions across all exo views. After filtering, we retain 110 hours of smooth and accurate EE4D-Motion data for UniEgoMotion training.